A person stands 6.00 m from a speaker, and 8.00 m from an identical speaker. What is the wavelength of the first (n=1) interference maximum (constructive & destructive)?

A person stands 6.00 m from a speaker, and 8.00 m from an identical speaker. What is the wavelength of the first (n=1) interference maximum (constructive)? Speed of sound 343 m/s) (Unit m) Enter
Answer #1

distance from speaker = r1 = 6 m

distance from identical speaker = r2 = 8m

for constructive interference path difference = r2 - r1 = n * lamda


                                                                       8-6 = 1* lamda

   wavelength = lamda = 2 m